Consider uniform electric field E=3*10^3 I ,calculate the flux of the field through a square of area 10 cm^2 when
1)its plane is parallel to y-z plane
2)the normal to its plane makes a 60° angle with the x axis

Grade:12

1 Answers

Khimraj
5 years ago
E = 3*103 A = 10*10-4,
i) \theta = 0
  then flux = EAcos\theta = 3 units
​ii) \theta = 60
  then flux = EAcos\theta = 3/2 units
Hope it clears. If you like anser then please approve it.